OpenStudy (anonymous):

For the following parabola equation, find the x-intercepts and the coordinates of the vertex. y= -x^2 -2x + 3

OpenStudy (anonymous):

are you familiar with completing the square?

OpenStudy (anonymous):

to find the x-intercept, simply let y=0.

OpenStudy (anonymous):

No I dont know completing the squares I've seen it but I could never get it.

OpenStudy (anonymous):

Okay, then are you familiar with differentiation?

OpenStudy (anonymous):

like... factoring?

OpenStudy (anonymous):

um, dy/dx. If not, then I guess you'll have to start from completing the square.

OpenStudy (anonymous):

ya lets start from completing the square idk what dy/dx is.

OpenStudy (anonymous):

\((x+1)^2 = x^2 +2x +1\) \((x+1)^2+3 = x^2 +2x +4\) So, completing the square is simply reverting the right side back to the left side. Basically , we try to guess what number fits the square \(y= -x^2 -2x + 3 \) now, take out the neg so it's easier. \(y= -(x^2 +2x - 3) \) Notice that x^2 +2x just almosts fits a perfect square. so, we add 1 to it. but to compensate, we minus 1 too, so it isn't like we're pulling it out of thin air. \(y= -(x^2 +2x+1-1 - 3) \) \(y= -(x^2 +2x+1 - 4) \) Now, \(x^2 +2x+1=(x+1)^2\) \(y= -( (x+1)^2- 4) \) \(y= - (x+1)^2+4 \) After completing the square, you can see that the max value for y will be when \((x+1)^2=0\) so, x=-1, y=4, vertex is (-1,4) you follow?

OpenStudy (anonymous):

woah, lets see

OpenStudy (anonymous):

ya I dont get how you get from y=−(x2+2x+1−4) to y=−((x+1)^2 −4) because with (x^2 + 2x +1 -4) you should get x^2 +2x -3

OpenStudy (anonymous):

yes. but in completing the square, we want to make an quad expression into a perfect square, with some thing behind. (x^2 + 2x +1 -4) looks very much like ((x+1)^2 −4)

OpenStudy (anonymous):

I dont think im getting it and my mind is getting flustered, I need a break.. Thank you for trying anyway.

OpenStudy (anonymous):

lol okaay